If the first and third of three consecutive odd integers are added, the result is 57 less than five times the second integer. Find the third integer.

Answer by Mathtut(3670)

ok lets call our 3 consecutive odd integers a+1, a+3,and a+5
so (a+1)+(a+5)=5(a+3)-57 solve for a..... 3a=48 so a=16 so first integer is a+1
which is 17 a+3 is 19 and a+5 which is what we are after is 21
answer:21